the length of a rectangle is fixed at 14 cm. what widths will make the perimeter greater than 70 cm.
Answer by Gogonati(809)

We know that 2L+2W=P or L+W=P/2=35 cm. In our problem L=14 cm and W=x cm.
Write the inequality:14+x≥35, solving this inequality we get : x≥21.
That this problem to be clear you need to interchange the dimensions: W= 14cm.